OpenStudy (anonymous):

sqrt (18x^7 over y^7)

OpenStudy (anonymous):

\[\sqrt{18 x^7/y^7}\]

OpenStudy (anonymous):

yep, that's it

OpenStudy (anonymous):

rationalize the denominator in the expression. Assume all variable represent positive real numbers.

OpenStudy (anonymous):

\[\sqrt{18x^7}/\sqrt{y^7}\times \sqrt{y^7}/\sqrt{y^7}\] \[\sqrt{18x^7y^7} \over y^7\]

OpenStudy (anonymous):

not it...

jimthompson5910 (jim_thompson5910):

You can simplify further \[\frac{\sqrt{18x^7y^7}}{y^7}\] \[\frac{\sqrt{18}\sqrt{x^7}\sqrt{y^7}}{y^7}\] \[\frac{3\sqrt{2}x^3\sqrt{x}y^3\sqrt{y}}{y^7}\] \[\frac{3x^3y^3\sqrt{2xy}}{y^7}\]

jimthompson5910 (jim_thompson5910):

Oops not done simplifying, you can simplify it even further to get \[\frac{3x^3\sqrt{2xy}}{y^4}\]

OpenStudy (anonymous):

you rock